# P3P Made Easy

### v1

@@ Abstract
We present a novel algebraic solution to the Perspective-Three-Point (P3P) problem, which aims to recover the absolute pose of a calibrated camera from three 2D–3D correspondences. Our method reformulates the problem into a quartic polynomial with coefficients that are analytically simple and computationally efficient. Despite its simplicity, the proposed solver achieves accuracy and runtime performance comparable to state-of-the-art methods. Extensive experiments on synthetic datasets validate its robustness and efficiency. This combination of simplicity and performance makes our solver appealing for both real-time systems and educational contexts, where interpretability and reliability are critical.
@@ 1 Introduction
[…]
Over the years, various P3P methods have been proposed, and they are typically classified by the degree and form of the polynomial equations they produce — most commonly quartic or cubic. While quartic-based solvers have been traditionally favored for their analytical simplicity, cubic-based solvers have recently attracted increasing interest due to their numerical stability and computational efficiency [17, 4].
In this work, we propose a novel quartic-based P3P solver that is numerically stable and computationally efficient. Our main contribution is twofold:
@@
    We show that the P3P problem can be reduced to solving a quartic polynomial with coefficients that are remarkably straightforward to derive and interpret. The derivation relies only on elementary mathematics, making the approach accessible even at the high school level. To the best of our knowledge, this degree of simplicity has not been previously demonstrated in the literature on the P3P problem.
@@
While many existing P3P solvers derive quartic or cubic polynomials using complex symbolic or geometric tools, our method stands out for its exceptional simplicity. We arrive at the quartic using only elementary algebraic reasoning, without relying on coordinate reparameterization [12, 15] or intermediate coordinate systems [2, 16]. This leads to a solver that is not only efficient and numerically stable but also easy to derive, implement, and verify, making it particularly well-suited for both practical deployment and instructional purposes.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 formally defines the P3P problem, and Section 4 introduces the proposed method. Evaluation results and analysis are presented in Sections 5 and 6, respectively. Finally, Section 7 concludes the paper.
@@ 3 Problem Definition
4 Proposed Method
4.1 Theory
@@
Substituting this into (17), we obtain an equation involving only $x^{2}$, $x$ and $y$, which we then simplify to isolate $y$ as a rational function of $x$:
@@ 5 Results
    Generate a random ground-truth rotation matrix $\mathbf{R}_{\text{gt}}$ and a translation vector of length 111, $\mathbf{t}_{\text{gt}}$.
@@
    Unique solutions: Solutions that have a valid rotation matrix [^5] We consider a rotation matrix to be valid if it meets the following conditions: (1) $|\det{(\mathbf{R})-1}|<10^{-6}$, (2) $\lVert\mathbf{R}^{\top}\mathbf{R}-\mathbf{I}\rVert_{L1}<10^{-6}$, and (3) $|1-\text{Quaternion norm of }\mathbf{R}|<10^{-5}$. and a reprojection error in the normalized image coordinates below $10^{-4}$, after removing the duplicates.
@@
The main results are presented in Tab. 1. It can be seen that Ding_new [4] is the most robust method. In terms of the number of ground-truth solutions obtained, our method is the second best: The difference between ours and Ding_new [4] is only 888 out of $10^{8}$ problems, amounting to a statistically negligible difference of 0.000008% of the total problems. In terms of the number of good solutions, the difference is only 4 out of $10^{8}$, which amounts to 0.000004%. We also found that the method by Wu et al. [23] performs slightly worse than the authors reported: In [23], it achieves the ground truth in 99.99997% of cases, while in our evaluation, it does so in 99.99989% of cases. We speculate that two potential causes may account for this discrepancy. First, we use a sample size that is 10 times larger than that used in [23], so there is a lower risk of sampling bias. Second, the authors of [23] might have mistakenly simulated the depths $d_{i}$ in (38) using a uniform distribution within the interval [0.1,100][0.1,100][ 0.1 , 100 ], instead of [0.1,10][0.1,10][ 0.1 , 10 ] as stated in their paper [^6] We suspect this for two main reasons: First, the authors’ public source code uses the distribution $\mathcal{U}(0.1,100)$ for the depths, and the file containing this line is identical to one uploaded by the authors of [4] where the same distribution is used. Second, our experiment confirmed that using the distribution $\mathcal{U}(0.1,100)$ for the depths improves the results of their method, producing numbers very close to those reported in [23]..
@@
Method Mean Median Maximum
Ours 1.514e-12 3.558e-14 9.662e-7
Wu [23] 6.918e-12 3.266e-14 9.666e-7
Ding_new [4] 1.335e-12 2.266e-14 8.306e-7
Ding_old [4] 4.647e-12 3.010e-14 9.925e-7
Persson [17] 1.082e-11 4.216e-14 9.960e-7
Ke [11] 9.195e-11 1.835e-14 1.000e-6
Kneip [12] 2.756e-10 4.317e-14 1.000e-6
Nakano [16] 1.694e-12 2.526e-14 9.589e-7
Time (ns)
Ours
Wu [23]
Ding_new [4]
Ding_old [4]
Persson [17]
Ke [11]
Kneip [12]
Nakano [16]
@@ 6 Discussion
A key strength of our method lies in its simplicity. By reducing the P3P problem to solving a single quartic polynomial with straightforward coefficients, we avoid the complexity often seen in other approaches. Despite its simplicity, our method achieves accuracy and runtime on par with state-of-the-art solvers. In many cases, it even outperforms more elaborate methods in terms of stability and robustness. This demonstrates that a well-designed minimal formulation can match or even outperform more sophisticated alternatives.
@@ 7 Conclusion
In this work, we revisited the P3P problem and showed that it can be solved by finding the roots of a quartic polynomial with compact, analytically simple coefficients. We also introduced several implementation strategies that significantly improve numerical stability in practice. Extensive evaluation demonstrated that our method delivers accuracy and runtime performance on par with state-of-the-art approaches. Beyond its empirical performance, the algebraic simplicity of our method makes it a valuable tool for teaching geometric vision concepts, developing lightweight systems for edge computing, and providing a clear baseline for future theoretical work on minimal solvers. Our code is publicly available at https://github.com/sunghoon031/P3P.

### v2

@@ Abstract
We revisit the classical Perspective-Three-Point (P3P) problem, which aims to recover the absolute pose of a calibrated camera from three 2D–3D correspondences. It has long been known that P3P can be reduced to a quartic polynomial with analytically simple and computationally efficient coefficients. However, this elegant formulation has been largely overlooked in modern literature. Building on the theoretical foundation that traces back to Grunert’s work in 1841, we propose a compact algebraic solver that achieves accuracy and runtime comparable to state-of-the-art methods. Our results show that this classical formulation remains highly competitive when implemented with modern insights, offering an excellent balance between simplicity, efficiency, and accuracy.
@@ 1 Introduction
P3P can also be viewed as a special case of the more general Perspective-$n$-Point (PnP) problem, which estimates camera pose from $n\geq 3$ 2D-3D correspondences. As the minimal case of PnP, P3P plays an essential role in a RANSAC-based frameworks [6], enabling the robust generation of pose hypotheses from minimal data. P3P has been a core component in numerous robust estimation pipelines, where its efficiency and stability directly impact the performance of the overall system.
Over the decades, various P3P methods have been proposed, and they are typically classified by the degree and form of the polynomial equations they produce — most commonly quartic or cubic. While quartic-based solvers have been traditionally favored for their analytical simplicity, cubic-based solvers have recently attracted increasing interest due to their numerical stability and computational efficiency [17, 4].
In this work, we propose a quartic-based P3P solver that is numerically stable and computationally efficient. Our main contribution is twofold:
@@
    We revisit the formulation of P3P and show that it reduces to a quartic polynomial with coefficients that are remarkably simple to derive and compute. This quartic was previously reported by Smith in 1965 [21], while an equivalent form, differing only by a scale factor, can be traced back to Grunert’s seminal work in 1841 [9].
@@
While many recent P3P solvers derive quartic or cubic polynomials using complex symbolic or geometric tools, our method stands out for its exceptional simplicity. It can be derived and understood using only elementary algebraic reasoning, as it does not rely on advanced techniques, such as coordinate reparameterization [12, 15] or intermediate coordinate systems [2, 16]. This leads to a solver that is not only efficient and numerically stable but also easy to derive, implement, and verify, making it particularly well-suited for both practical deployment and instructional purposes.
@@ 3 Problem Definition
4 Method
4.1 Theory revisited
@@
Substituting this into (17), we obtain an equation involving only $^{2}, xx$ and $y$, which we then simplify to isolate $y$ as a rational function of $x$:
@@
We note that the quartic formulation (26)–(31) was previously derived and used by Smith in 1965 [21]. A mathematically equivalent form, differing only by a scale factor, was originally derived by Grunert in 1841 [9]. Another scaled, yet equivalent, variant was proposed by Fischler and Bolles in 1981 [6]. Table 1 compares the quartics employed in these works. Although all these quartic formulations are mathematically equivalent, we found that ours (identical to Smith’s [21]) yields the highest numerical precision.
Method Scale factor relative to ours (26)
Grunert (1841) [9] $\left(s_{13}\right)^{-2}$
Smith (1965) [21] 1-1
Fischler & Bolles
(1981) [6]
$\left(\frac{s_{12}}{s_{13}s_{23}}\right)^{2}$
Table 1: The relative scale factors between our quartic (26) and those presented in earlier works [9, 21, 6].
@@ 5 Results
    Generate a random ground-truth rotation matrix $\mathbf{R}_{\text{gt}}$ and a translation vector of length $1$, $\mathbf{t}_{\text{gt}}$.
@@
[…]
@@
The main results are presented in Tab. 2. It can be seen that Ding_new [4] is the most robust method. In terms of the number of ground-truth solutions obtained, our method is the second best: The difference between ours and Ding_new [4] is only $8$ out of $10^{8}$ problems, amounting to a statistically negligible difference of 0.000008% of the total problems. In terms of the number of good solutions, the difference is only 4 out of $10^{8}$, which amounts to 0.000004%. We also found that the method by Wu et al. [24] performs slightly worse than the authors reported: In [24], it achieves the ground truth in 99.99997% of cases, while in our evaluation, it does so in 99.99989% of cases. We speculate that two potential causes may account for this discrepancy. First, we use a sample size that is 10 times larger than that used in [24], so there is a lower risk of sampling bias. Second, the authors of [24] might have mistakenly simulated the depths $d_{i}$ in (38) using a uniform distribution within the interval $[0.1,100]$, instead of $[0.1,10]$ as stated in their paper [^6] This suspicion has not yet been confirmed by the authors of [24], as the direct correspondence with them may potentially violate the anonymity policy for the reviewing process. However, we have two reasons why we suspect that this might be the case. First, the authors’ public source code uses the distribution $\mathcal{U}(0.1,100)$ for the depths, and the file containing this line is identical to one uploaded by the authors of [4] where the same distribution is used. Second, our experiment confirmed that using the distribution $\mathcal{U}(0.1,100)$ for the depths improves the results of their method, producing numbers very close to those reported in [24]..
@@
Time (ns) Ours Wu [24] Ding_new [4] Ding_old [4] Persson [17] Ke [11] Kneip [12] Nakano [16]
@@ 6 Discussion
By reducing the P3P problem to solving a single quartic polynomial with straightforward coefficients, we avoid the complexity often seen in other approaches. Despite its simplicity, our method achieves accuracy and runtime on par with state-of-the-art solvers. In many cases, it even outperforms more elaborate methods in terms of stability and robustness. This demonstrates that a well-designed minimal formulation can match or even outperform more sophisticated alternatives.
@@ 7 Conclusion
In this work, we revisited the classical P3P problem and demonstrated that it can be solved efficiently by finding the roots of a quartic polynomial with compact, analytically simple coefficients. This quartic formulation, originally presented by Smith in 1965 [21], is mathematically equivalent to those of Grunert (1841) [9] and Fischler and Bolles (1981) [6], differing only by scale factors. Several implementation strategies were introduced to enhance numerical stability, leading to accuracy and runtime performance comparable to state-of-the-art methods.
Beyond its empirical results, the algebraic simplicity of our formulation makes it well suited for teaching geometric vision concepts, deploying lightweight systems on edge devices, and serving as a clear baseline for future research on minimal solvers. Our work also illustrates that classical formulations, when revisited with modern insight, can rival or even surpass many recent, more elaborate methods. Our code is publicly available at https://github.com/sunghoon031/P3P.
Method Mean Median Maximum
Ours 1.514e-12 3.558e-14 9.662e-7
Wu [24] 6.918e-12 3.266e-14 9.666e-7
Ding_new [4] 1.335e-12 2.266e-14 8.306e-7
Ding_old [4] 4.647e-12 3.010e-14 9.925e-7
Persson [17] 1.082e-11 4.216e-14 9.960e-7
Ke [11] 9.195e-11 1.835e-14 1.000e-6
Kneip [12] 2.756e-10 4.317e-14 1.000e-6
Nakano [16] 1.694e-12 2.526e-14 9.589e-7